Can someone help solving this problem!!
Soybean meals is 18% protein; cornmeal is 9% protein. How many pounds of each should be mixed together in order to get 360-1b mixture that is 11% protein.
How many pounds of the corn meal should be in the mixture?_____

Answer by mananth(16949)

Soybean meals is 18% protein; cornmeal is 9% protein. How many pounds of each should be mixed together in order to get 360-1b mixture that is 11% protein.
..
Soybean meals is 18% protein; : x lb
cornmeal is 9% protein. : 360-x lb
360-1b mixture that is 11% protein.
..
Concentration equation
18*x+9(360-x)=11*360
18x+3240-9x=3960
9x+3240=3960
-3240
9x=3960-3240
9x=720
/9
x= 80 lb soyabean
y = 360-x = 240 lb cornmeal
...
m.ananth@hotmail.ca